# Reducing the Scope of Language Models

### v2

@@ B.5 Accepting Multiple Tasks
For Section 4.3, we use the following accept/reject sets for each group of plots:
@@ B.6 Robustness to Adversarial Prompts
For Section A.3, we use the following accept/reject sets for each group of plots:
@@
  • Depth: 2 (Mehrotra et al. (2023) point out that a shallower, wider tree is best)
@@ B.7 Precise Scoping
As described in Section A.4, the finegrained accept (FA) set is a single task taken from the same category as the experiment (SA, S, PE respectively). We then make sure that the finegrained reject (FR) set does not contain the dataset that task was drawn from, and allow for all other tasks in the category.
@@ B.8 Effect of Data Quantity
For Section A.6, we use the following accept/reject sets for each group of plots:
@@ B.9 Effect of LoRA Rank
For Section A.7, we use the following accept/reject sets for each group of plots:
@@ B.10 Representation Analysis
In Section 4.4, for each instance, we compute average cosine similarity for every token at every layer between the base model (Sys.) and the model we are evaluating. Then we truncate these matrices to the minimum instance length in the evaluation set, and average all cosine similarities over instances. Thus we are looking at the average tail representations of instances in the dataset.
@@ B.11 Code Sources
We leverage PyTorch (Paszke et al., 2019) for training, HuggingFace Transformers (Wolf et al., 2020) for models and training infrastructure, HuggingFace Datasets (Lhoest et al., 2021) for particular datasets, and HuggingFace TRL (von Werra et al., 2020) for DPO. We also work on top of the official TAP (Mehrotra et al., 2023) for our adversarial evaluations, and found the source code for CB (Zou et al., 2024) quite helpful for verifying our implementation. We aggregate results with NumPy (Harris et al., 2020) and Pandas (McKinney et al., 2011) and make plots with Matplotlib (Hunter, 2007).
@@ Appendix C Sample outputs
To further justify the design of our detectors, we provide some sample outputs in Table 3. Notice how CB produces repetitive tokens, we find this pattern quite common on manual inspection, thus we base our evaluation on detecting such repetitions.

### v3

@@ B.5 Accepting Multiple Tasks
For Section of Accepting multiple tasks, we use the following accept/reject sets for each group of plots:
@@ B.6 Robustness to Adversarial Prompts
For Section A.4 of Robustness to Adversarial Prompts, we use the following accept/reject sets for each group of plots:
@@
  • Depth: 2 (mehrotra2023tree point out that a shallower, wider tree is best)
@@ B.7 Precise Scoping
As described in Section A.5 of Precise Scoping, the finegrained accept (FA) set is a single task taken from the same category as the experiment (SA, S, PE respectively). We then make sure that the finegrained reject (FR) set does not contain the dataset that task was drawn from, and allow for all other tasks in the category.
@@ B.8 Effect of Data Quantity
For Section A.7 of Effect of Data Quantity, we use the following accept/reject sets for each group of plots:
@@ B.9 Effect of LoRA Rank
For Section A.8 of Effect of LoRA Rank, we use the following accept/reject sets for each group of plots:
@@ B.10 Representation Analysis
In Section of Representation Analysis, for each instance, we compute average cosine similarity for every token at every layer between the base model (Sys.) and the model we are evaluating. Then we truncate these matrices to the minimum instance length in the evaluation set, and average all cosine similarities over instances. Thus we are looking at the average tail representations of instances in the dataset.
@@ B.11 Code Sources
We leverage PyTorch (paszke2019pytorch) for training, HuggingFace Transformers (wolf2020transformers) for models and training infrastructure, HuggingFace Datasets (lhoest2021datasets) for particular datasets, and HuggingFace TRL (vonwerra2022trl) for DPO. We also work on top of the official TAP (mehrotra2023tree) for our adversarial evaluations, and found the source code for CB (zou2024improving) quite helpful for verifying our implementation. We aggregate results with NumPy (harris2020array) and Pandas (mckinney2011pandas) and make plots with Matplotlib (Hunter:2007).
@@ Appendix C Sample outputs
To further justify the design of our detectors, we provide some sample outputs in Table 4 of Sample rejection outputs. Notice how CB produces repetitive tokens, we find this pattern quite common on manual inspection, thus we base our evaluation on detecting such repetitions.